OpenStudy (anonymous):

Can someone please help me? Find all solutions in the interval [0, 2π). (sin x)(cos x) = 0

OpenStudy (anonymous):

either \(\cos(x)=0\) or \(\sin(x)=0\) do you know where?

OpenStudy (anonymous):

I really don't know how to do this at all.

OpenStudy (anonymous):

should have these memorized. \(\sin(0)=0,\sin(\pi)=0,\sin(2\pi)=0,...\)

OpenStudy (anonymous):

as for cosine, \(\cos(\frac{\pi}{2})=0,\cos(\frac{3\pi}{2})=0,...\)

OpenStudy (anonymous):

look at the picture of the unit circle on the attached sheet sheet. sine is the first coordinate, cosine is the second, see for which angles the first coordinate is zero (for cCED12: Can someone please help me? Find all solutions in the interval [0, 2π). … CED12: Can someone please help me? Find all solutions in the interval [0, 2π). … osine) or the second coordinate is zero (for sine)

OpenStudy (anonymous):

Ok so out of the following possible answers pi/2, pi 0, pi/2, pi, 3pi/2 pi, 3pi/2 0, 3pi/2 the right answer is pi/2, pi right?